# PAPER: Privacy-Preserving ResNet Models using Low-Degree Polynomial Approximations and Structural Optimizations on Leveled FHE

### v1

@@ Appendix C Why the Penalty Function Works?
The first term is the standard cross‑entropy loss computed from the network output $g(x)$. The second term is a layer-specific penalty that acts only on the pre-activations at layer $l$, penalizing values that fall outside the clipping interval. A gradient-descent step with learning rate $\eta$ updates the weight matrix by
@@
Putting it all together, we obtain the total gradient-based update for $W^{(l)}$
@@
$\Delta z_{p}^{(l)}$ $\;=\;\overline{z}_{p}^{(l)}\;-\;z_{p}^{(l)}$
$\;=\;\overline{W}^{(l)}\;h^{(l-1)}_{p_{d}}(x)\;-\;W^{(l)}\;h^{(l-1)}_{p_{d}}(x)$
@@
we obtain the penalty components $\Delta z_{\rm CE}^{(l)}$ and $\Delta z_{\rm pen}^{(l)}$ as mentioned in the lemma. ∎
@@
whenever $\left\|d^{(l)}\right\|_{2}\neq 0$. Therefore, $\Delta z_{\mathrm{pen}}^{(l)}$ “pulls back” each element of $z_{p}^{(l)}$ lying outside [c,c][-c,c] towards the clipping interval.
The clipping residual $d^{(l)}$ is defined as the difference between the current pre‑activation $z_{p}^{(l)}$ and its clipped version. $\Delta z_{\rm pen}^{(l)}$ is the component in $\Delta z_{p}^{(l)}$ induced by the clip-range penalty. Taking the inner product of $\Delta z_{\rm pen}^{(l)}$ with the clipping residual $d^{(l)}$ gives
@@
$\;=\;-\eta\;\zeta\;\left\|h^{(l-1)}_{p_{d}}(x)\right\|_{2}^{2}\;\frac{1}{\left\|d^{(l)}\right\|_{2}}\;\langle d^{(l)},\,d^{(l)}\rangle$
@@
Each factor on the right-hand side: $\eta$ (learning rate), $\zeta$ (regularization parameter), $\left\|h^{(l-1)}_{p_{d}}(x)\right\|_{2}^{2}$ (squared norm of the preceding hidden representation) and $\left\|d^{(l)}\right\|_{2}$ (magnitude of the clipping residual) is strictly positive whenever any pre‑activation lies outside [c,c][-c,c]. Hence the entire product is strictly negative
@@
Appendix D Training Stability during Polynomial Training
D.1 Pre-Activation Clipping
During early training stages, network weights remain close to their random initialization, and the optimization process does not immediately constrain the pre-activation values to lie within the target approximation interval [c,c][-c,c]. As a result, unbounded pre-activations may arise before the model learns to keep them within range, leading to training instability that can degrade model behavior beyond recovery. To prevent this, we restrict pre-activations to the interval [c,c][-c,c] before evaluating the polynomial activation, which is achieved through a clipping strategy $\mathrm{clip}(z;[-c,c])$ (Diaa et al., 2024):
Importantly, this clipping operation is applied only during training to stabilize learning. It is performed after computing the clip-range penalty to ensure that gradients from the regularization term, which encourages the network to keep pre-activations within the approximation interval [c,c][-c,c], are preserved and not masked by the clipping. At inference time, the clipping function is removed:
$h^{(l)}_{p_{d}}(x)\;=\;p_{d}\left(W^{(l)}h^{(l-1)}_{p_{d}}(x)\right).$
The model, having learned to constrain pre-activations during training, is expected to remain within the approximation interval without explicit clipping at inference.
D.2 Regularization Warm-up
While activation regularization and pre-activation clipping are both necessary for training stability, applying the full regularization strength from the outset can lead to numerical instability, particularly in larger models. In the early epochs, many pre-activation values lie outside the target interval [c,c][-c,c] across several layers. The clip-range penalty adds a contribution from every such layer, so the penalty term becomes very large. In extreme cases, this can cause the total loss to become numerically undefined.
To address this issue, the regularization strength $\zeta$ is progressively increased over the initial training epochs, which is implemented through a regularization warm up schedule (Diaa et al., 2024). Let $T_{warm}$ denote the total number of warm-up epochs. For each epoch $t$, we define a time-dependent regularization weight $\zeta_{t}$ as:
where $\{\alpha_{t}\}_{t=1}^{T_{warm}}$ is a fixed sequence of scaling factors satisfying $0<\alpha_{1}<\cdots<\alpha_{T_{warm}}<1$. In practice, we empirically find that setting $T_{warm}=4$ provides a stable convergence. We use a simple predefined sequence of scaling factors: $\alpha=\left\{\frac{1}{100},\frac{1}{50},\frac{1}{10},\frac{1}{5}\right\}$, which produces the epoch-wise regularization strength: $\zeta_{t}\;\in\;\left\{\frac{\zeta}{100},\frac{\zeta}{50},\frac{\zeta}{10},\frac{\zeta}{5},\zeta,\zeta,\dots\right\}$. This warm-up schedule ensures that the regularization penalty is introduced progressively in the initial training epochs without suffering from exploding loss values.
Appendix E Additional Details on Structural Optimizations
E.1 Detailed Derivations of Node Fusing
@@
$\quad+\underbrace{\big[b_{X1}\big(2c_{2}(b_{X0}+b_{Y0})+c_{1}\big)\big]}_{d_{X}}x+\underbrace{\big[b_{Y1}\big(2c_{2}(b_{X0}+b_{Y0})+c_{1}\big)\big]}_{d_{Y}}y$
@@
E.2 Detailed Derivations of Weight Redistribution
E.2.1 Update Forward
@@
$\bar{c}_{ij}$ $=c_{ij}\upsilon^{-1}\quad\forall i\in\{0,\dots,d-1\},\forall j\in\{0,\dots,d-i\}.$
@@
$\bar{c}_{ij}$ $=c_{ij}\upsilon^{i}\quad\forall i\in\{0,\dots,d\},\forall j\in\{0,\dots,d-i\}.$
E.2.2 Update Backward
@@
$\bar{c}_{ij}$ $=c_{ij}\upsilon^{-i}\quad\forall i\in\{0,\dots,d-1\},\forall j\in\{0,\dots,d-i\}.$
@@
$\bar{c}_{ij}$ $=c_{ij}\upsilon\quad\forall i\in\{0,\dots,d\},\forall j\in\{0,\dots,d-i\}.$
E.3 Example of Tower Reuse
Suppose following an activation function, there is a convolution. Let modulus $q_{i}\approx\Delta^{2}$, thus $\lambda(q_{i})=2$. Consider an input $x$ with scaling factor $\Delta_{x}=\Delta$, giving $\lambda(x)=1$. Applying a quadratic activation $y=x^{2}+c_{1}x+c_{0}$ yields $\lambda(y)=2$, with $\lambda(c_{1})=1$ and $\lambda(c_{0})=2$. Next, a convolution $z=\sum_{i}\omega_{i}y_{i}+\alpha$ is performed. Since $\lambda(y)=2$, by setting $\lambda(\omega)=1$, we obtain $\lambda(z)=3$. As $\lambda(z)>\lambda(q_{i})$, a rescale is triggered. The rescale operation drops $q_{i}$ from the modulus chain, reducing the level $\Lambda(\cdot)$ by one and resetting the sublevel to one:
E.4 Level Analysis
Table 1: Number of levels required for multiplications in ResNet models. $P_{4}$: Degree-4 polynomial activation used in PILLAR (Diaa et al., 2024). $P_{2}$: Degree-2 polynomial activation (§2). $P_{2}F$: $P_{2}$ with node fusing (§3.1). $P_{2}R$: $P_{2}$ with weight redistribution (§3.2). $P_{2}FR$: $P_{2}$ with node fusing and weight redistribution. $P_{2}FRT$: $P_{2}FR$ with tower reuse (§3.3).
Model $P_{4}$ $P_{2}$ $P_{2}F$ $P_{2}R$ $P_{2}FR$ $P_{2}FRT$
ResNet-18 87 70 53 35 35 18
ResNet-20 97 78 59 39 39 20
ResNet-32 157 126 95 63 63 32
Table 1 summarizes the number of levels $\mathcal{L}$ required for multiplication across different ResNet variants under various optimization techniques. To illustrate, we focus on ResNet-18, which comprises 17 convolutional layers, batch normalization layers, and activation functions, along with a single average pooling and linear layer:
  • $P_{4}$: Prior work achieving the lowest $\mathcal{L}$ (Diaa et al., 2024) uses a degree-4 polynomial as activation function. This results in $\mathcal{L}=87$, with each convolution, batch normalization, average pooling, and linear layer contributing $\mathcal{L}=1$, and the polynomial activation $\mathcal{L}=3$.
  • $P_{2}$: By replacing the activation function with a degree-2 polynomial, as described in §2, the required level for activations drops to 2, reducing the model $\mathcal{L}$ to 70.
  • $P_{2}F$: Applying node fusing (§3.1) to $P_{2}$ removes all batch normalizations, lowering $\mathcal{L}$ to 53.
  • $P_{2}R$: Alternatively, applying weight redistribution (§3.2) to $P_{2}$ yields $\mathcal{L}=35$. This is achieved because the highest-order coefficients in batch normalization and polynomial activations are normalized to one, simplifying them to $x+b_{0}$ ($\mathcal{L}=0$) and $x^{2}+c_{1}x+c_{0}$ ($\mathcal{L}=1$), respectively. Moreover, the divisor in average pooling is also normalized to one, giving it $\mathcal{L}=0$. These simplifications collectively result in $\mathcal{L}=35$.
  • $P_{2}FR$: Combining node fusing and weight redistribution retains $\mathcal{L}=35$. However, it requires fewer homomorphic operations overall, since node fusing eliminates batch normalizations entirely.
  • $P_{2}FRT$: Finally, applying tower reuse (§3.3) to $P_{2}FR$ allows a non-linear polynomial activation followed by a linear kernel function to share the same level. This halves the total levels to $\mathcal{L}=18$, following $\mathcal{L}_{P_{2}FRT}=\lceil\frac{\mathcal{L}_{P_{2}FR}}{2}\rceil$ for ResNet models.
In summary, our techniques reduce levels by nearly a factor of five compared to prior work.
Appendix F Illustration of Slice-wise Clustering
F.1 Single Model Slice Clustering
Figure 3 illustrates how convolution filters are divided into slices along the width axis $W_{l}$, with each colored slice $S^{(l)}_{s}$ clustered independently using its own codebook. Each 3D block in the figure corresponds to a convolution filter associated with one output channel. The three axes correspond to kernel height $H_{l}$ (vertical), kernel width $W_{l}$ (horizontal), and input channels $I_{l}$ (depth). The collection of such filters forms the output channels $O_{l}$. Slices are taken as vertical strips along the width axis $W_{l}$, and the figure highlights them with distinct colors. The blue strip corresponds to slice $S^{(l)}_{1}$, the yellow strip to slice $S^{(l)}_{2}$, and the pink strip to slice $S^{(l)}_{3}$. Each slice $S^{(l)}_{s}$ groups together all weights $W^{(l)}_{o,i,h,s}$} across output channels $O_{l}$, input channels $I_{l}$, and kernel height $H_{l}$ at a fixed width position $s$, and is clustered independently with its own codebook $\mathcal{C}^{(l)}_{s}$. This illustration emphasizes that clustering is performed slice by slice, rather than across the entire filter volume.
Figure 3: Illustration of single model slice-wise clustering.
F.2 Ensemble Slice Clustering
Figure 4 illustrates how slice-wise clustering is extended from a single model to an ensemble of polynomial networks.
Figure 4: Illustration of ensemble slice-wise clustering.
Each 3D block corresponds to the convolution filters of one model instance, spanning kernel height $H_{l}$ (vertical), kernel width $W_{l}$ (horizontal), and input channels $I_{l}$ (depth). The collection of such filters across output channels $O_{l}$ forms the layer weights $W^{(l,m)}$ for model $m$. For a fixed layer $l$, slices are again taken as vertical strips along the width axis $W_{l}$, with colors indicating different slices $S^{(l,m)}_{s}$ within each model. Unlike the single-model case, however, we now align slices from multiple models at the same width position $s$ and stack them together. This produces a slice matrix $X^{(l)}_{s}\in\mathbb{R}^{N_{s}\times M}$, where each row corresponds to one weight coordinate across all $M$ models, and $N_{s}=O_{l}I_{l}H_{l}$. Clustering is performed jointly in $\mathbb{R}^{M}$, producing a shared codebook $\mathcal{C}^{(l)}_{s}$ for slice $s$. The illustration emphasizes that ensemble clustering is not applied to each model independently. Instead, slices are stacked across models at the same spatial position, and clustering is performed jointly.
Appendix G Ablation Study
G.1 Ablation on Clipping Range (Value of $c$)
We conduct an ablation study to evaluate the effect of the clipping interval [c,c][-c,c] on model performance. The choice of $c$ determines how much of the activation distribution is preserved versus truncated: if $c$ is too small, informative activations are excessively clipped, which can harm representational capacity, while if $c$ is too large, the clipping mechanism provides little stabilization benefit, allowing unstable activations to propagate. Figure 5 shows the results for CIFAR-10 with ResNet-18.
Figure 5: Results of the clipping range ablation study on CIFAR-10 with ResNet-18. (a) Distribution of pre-activation inputs to ReLU across the entire network, showing concentration within [2,2][-2,2]. (b) Test accuracy under different clipping intervals, where [2,2][-2,2] produces the best performance.
Figure 5(a) presents the distribution of pre-activation inputs to ReLU across the entire network, where the values are concentrated within the interval [2,2][-2,2] and only a small fraction lies outside this range, supporting the choice of restricting pre-activations to [2,2][-2,2]. Figure 5(b) reports the accuracy obtained under different clipping intervals. While several ranges were tested, [2,2][-2,2] consistently achieved the highest mean accuracy across runs. Smaller intervals degraded performance by removing too many activations, whereas larger intervals reduced stability and resulted in lower accuracy. The same trend holds across other dataset–architecture combinations. Based on these observations, we adopt [2,2][-2,2] as the clipping range in all our experiments.
G.2 Ablation on Penalty Strength (Value of $\zeta$)
We conduct an ablation study to evaluate the effect of the penalty strength $\zeta$ on model performance. The parameter $\zeta$ controls the magnitude of the regularization term that penalizes pre-activations outside the target interval [c,c][-c,c]. Figure 6 shows the results for CIFAR-10 with ResNet-18.
Figure 6: Results of the penalty strength ablation study on CIFAR-10 with ResNet-18, showing test accuracy across different values of $\zeta$. The best performance is achieved at $\zeta=10^{-3}$; smaller values fail to regularize effectively, while larger values start to decrease accuracy.
The highest accuracy is obtained for $\zeta=10^{-3}$. Values smaller than $10^{-3}$ do not provide sufficient regularization and result in poor accuracy, while values greater than $10^{-3}$ start decreasing performance, indicating that excessive penalization constrains the model too strongly. The same trend holds across other dataset–architecture combinations. Based on these results, we set $\zeta=10^{-3}$ in all our experiments.
G.3 Comparison with PILLAR
We evaluate PILLAR using its open-source implementation (git, 2024b). PILLAR is designed for interactive PPML, which is a different use case from our non-interactive setting, and its repository does not provide code for private inference. For these reasons, we compare only in terms of accuracy and do not evaluate inference time. Table 2 reports accuracies (mean $\pm$ standard deviation) for ResNet-18, ResNet-20, and ResNet-32 on CIFAR-10 and CIFAR-100 using ReLU, PAPER (our method), and two versions of PILLAR. The original PILLAR is defined with a degree-4 polynomial. For a fair comparison, we additionally adapted this implementation to use a degree-2 polynomial, allowing a direct comparison to PAPER. On average, the adapted degree-2 version shows an accuracy drop of 6.5% with very high variance, particularly on CIFAR-100. The original degree-4 version is more stable but still falls short of ReLU by about 5.4% on average. In contrast, PAPER with degree-2 consistently stays much closer to ReLU, with only about a 1.3% drop on average. Importantly, PAPER remains stable across both datasets and all ResNet variants, avoiding the large fluctuations seen in PILLAR. These results demonstrate that while PAPER can exploit low-degree polynomials effectively, PILLAR either collapses when reduced to degree 2 or, in its intended degree-4 form, does not reach comparable accuracy.
Table 2: Classification accuracies (mean $\pm$ standard deviation) on CIFAR-10 and CIFAR-100 for ResNet-18, ResNet-20, and ResNet-32 using ReLU, PAPER (degree-2), and two versions of PILLAR. The original PILLAR method is degree-4 (Diaa et al., 2024), while the degree-2 variant is our adaptation of its open-source implementation for comparison.
Model Dataset ReLU PAPER PILLAR ($d=2$) PILLAR ($d=4$)
ResNet-18 CIFAR-10 94.57 $\pm$ 0.23 93.73 $\pm$ 0.27 91.78 $\pm$ 0.13 93.26 $\pm$ 0.15
CIFAR-100 76.22 $\pm$ 0.44 74.96 $\pm$ 0.51 71.92 $\pm$ 0.28 74.81 $\pm$ 0.22
ResNet-20 CIFAR-10 94.84 $\pm$ 0.25 94.06 $\pm$ 0.26 90.43 $\pm$ 0.40 90.98 $\pm$ 0.29
CIFAR-100 74.96 $\pm$ 0.36 73.89 $\pm$ 0.37 62.30 $\pm$ 3.96 62.55 $\pm$ 7.51
ResNet-32 CIFAR-10 95.23 $\pm$ 0.30 93.86 $\pm$ 0.24 90.40 $\pm$ 1.73 92.24 $\pm$ 0.27
CIFAR-100 76.58 $\pm$ 0.40 73.93 $\pm$ 0.53 66.38 $\pm$ 1.78 66.10 $\pm$ 1.12
Appendix H Additional Experimental Results
H.1 Additional Details of FHE Parameters
@@
For each model, we detail the encryption parameters used in our experiments. Our framework automatically determines the RNS moduli according to the specified sizes $|q_{i}|$. We chose the smallest values of $\Delta$ and $|q_{i}|$ that preserve accuracy during private inference and adjust $N$ for $\approx$128-bit security. The moduli are ordered as $q_{0},\dots,q_{L},P$, where $q_{0}$ holds the output, $q_{1},\ldots,q_{L}$ are for rescaling, and $P$ is for modulus expansion in relinearization. FHE terminology is provided in §A.

### v2

@@ Appendix A Why the Penalty Function Works?
The first term is the standard cross‑entropy loss from the network output $g(x)$. The second term is a layer-specific penalty that acts only on the pre-activations at layer $l$, penalizing values that fall outside the clipping interval. A gradient-descent step with learning rate $\eta$ updates the weight matrix by
@@
Putting it all together, we obtain the total update for $W^{(l)}$
@@
$\Delta z_{p}^{(l)}$ $\;=\;\overline{z}_{p}^{(l)}\;-\;z_{p}^{(l)}\;=\;\overline{W}^{(l)}\;h^{(l-1)}_{p_{d}}(x)\;-\;W^{(l)}\;h^{(l-1)}_{p_{d}}(x)$
@@
we obtain the penalty components $\Delta z_{\rm CE}^{(l)}$ and $\Delta z_{\rm pen}^{(l)}$. ∎
@@
whenever $\left\|d^{(l)}\right\|_{2}\neq 0$. Therefore, $\Delta z_{\mathrm{pen}}^{(l)}$ “pulls back" each element of $z_{p}^{(l)}$ lying outside [c,c][-c,c] towards the clipping interval.
The clipping residual $d^{(l)}$ is defined as the difference between the pre‑activation $z_{p}^{(l)}$ and its clipped version. $\Delta z_{\rm pen}^{(l)}$ is the component in $\Delta z_{p}^{(l)}$ induced by the clip-range penalty. Taking the inner product of $\Delta z_{\rm pen}^{(l)}$ with $d^{(l)}$ gives
@@
Each factor on the right-hand side: $\eta$ (learning rate), $\zeta$ (regularization parameter), $\left\|h^{(l-1)}_{p_{d}}(x)\right\|_{2}^{2}$ (squared norm of the preceding hidden representation) and $\left\|d^{(l)}\right\|_{2}$ (magnitude of the clipping residual) are strictly positive whenever any pre‑activation lies outside [c,c][-c,c]. Hence the entire product is strictly negative
@@
Appendix B Details on Structural Optimizations
B.1 Derivations of Node Fusing
@@
$\quad+\underbrace{\big[b_{X1}\big(2c_{2}(b_{X0}+b_{Y0})+c_{1}\big)\big]}_{d_{X}}x$
$\quad+\underbrace{\big[b_{Y1}\big(2c_{2}(b_{X0}+b_{Y0})+c_{1}\big)\big]}_{d_{Y}}y$
@@
B.2 Derivations of Weight Redistribution
B.2.1 Update Forward
@@
$\bar{c}_{ij}$ $=c_{ij}\upsilon^{-1}\qquad\begin{aligned} &\forall i\in\{0,\dots,d-1\},\\ &\forall j\in\{0,\dots,d-i\}.\end{aligned}$
@@
$\bar{c}_{ij}$ $=c_{ij}\upsilon^{i}\qquad\begin{aligned} &\forall i\in\{0,\dots,d\},\\ &\forall j\in\{0,\dots,d-i\}.\end{aligned}$
B.2.2 Update Backward
@@
$\bar{c}_{ij}$ $=c_{ij}\upsilon^{-i}\qquad\begin{aligned} &\forall i\in\{0,\dots,d-1\},\\ &\forall j\in\{0,\dots,d-i\}.\end{aligned}$
@@
$\bar{c}_{ij}$ $=c_{ij}\upsilon\qquad\begin{aligned} &\forall i\in\{0,\dots,d\},\\ &\forall j\in\{0,\dots,d-i\}.\end{aligned}$
Appendix C Details on FHE Parameters
@@
For each model, we detail the encryption parameters used in our experiments. Our framework automatically determines the RNS moduli according to the specified sizes $|q_{i}|$. We chose the smallest values of $\Delta$ and $|q_{i}|$ that preserve accuracy during private inference and adjust $N$ for $\approx$128-bit security. The moduli are ordered as $q_{0},\dots,q_{L},P$, where $q_{0}$ holds the output, $q_{1},\ldots,q_{L}$ are for rescaling, and $P$ is for modulus expansion in relinearization. FHE terminology is provided in §2.1.
VGG-16: $N=2^{15}$, $\Delta=2^{25}$, $\log_{2}{Q}=6\cdot 50+10\cdot 49+26+54=870$. Moduli: {0x4000000120001, 0x3ffffffdf0001, 0x40000001b0001, 0x3ffffffd20001, 0x4000000270001, 0x3ffffffcd0001, 0x20000000b0001, 0x20000001a0001, 0x1fffffff50001, 0x1ffffffea0001, 0x20000003b0001, 0x1ffffffd40001, 0x1ffffffba0001, 0x20000005e0001, 0x1ffffffb40001, 0x20000006d0001, 0x3ff0001, 0x3fffffffd60001}.